OpenStudy (anonymous):

Simplify the radical expression. sqrt(14q)*2sqrt(4q)

OpenStudy (anonymous):

\[\sqrt{14q}*2\sqrt{4q}\]

OpenStudy (anonymous):

whats the \[\sqrt{4}\]?

OpenStudy (anonymous):

what do you mean? this?\[\sqrt{2*2}\]

OpenStudy (anonymous):

yep. since \[\sqrt{4} = 2\] we pull out the two and now we have \[2*2 \sqrt{14q} *\sqrt{q}\] so now we want to multiply \[\sqrt{14q} * \sqrt{q} = \sqrt{14q^2}\] take half of the q power and more it out of the sqrt and now we \[q(2*2) \sqrt{14} = 4q \sqrt{14}\]

OpenStudy (anonymous):

omg, I got that and it looked wrong...I did not quite do it that way but I got the same answer...I took up way more paper, lol

OpenStudy (anonymous):

lol :) well as long as we got the same answer there is higher percentage that we are right!

OpenStudy (anonymous):

thanks for the assurance, at least I know I am now understanding some of what I am doing ;p

OpenStudy (anonymous):

:) well good luck with your course!